OpenStudy (anonymous):

if a perimeter of a rectangle is 72m & width is 16m then what is the area?

OpenStudy (anonymous):

area of a rectangle is width times length. so here, length =72 width=16 72 times 16 =1152

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

P = 2(L + W), P = 72, W = 16 72 = 2(L + 16) L + 16 = 36 L = 20 A = (20 cm)(16 cm) = 320 cm^2.